OpenStudy (anonymous):

A jar contains 7 blue cubes, 4 blue spheres, 5 green cubes, and 6 green spheres. If you select an object at random, what is the probability that the object is green or a cube? A. 23/22 B. 5/22 C. 9/11 D. 4

OpenStudy (anonymous):

B. 5/22

OpenStudy (anonymous):

thank you !

OpenStudy (blurbendy):

Add all the stuff up: 7 + 4 + 5 + 6 = 11 + 11 = 22 probability of a green object or a cube, so add the green objects and the total cubes 11 green objects + 12 cubes. = 23 23 / 22

OpenStudy (blurbendy):

no wait, i think i double counted

OpenStudy (anonymous):

Its B!

OpenStudy (blurbendy):

No, it's C! Here's why.

OpenStudy (blurbendy):

The only thing we're NOT counting are the BLUE spheres.

OpenStudy (blurbendy):

There are four of them, so 22 total objects - 4 blue spheres = 18 / 22 = 9 / 11

OpenStudy (anonymous):

Your right,

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

I am sooo sorry!

OpenStudy (anonymous):

No best response?